OpenStudy (anonymous):

what are the measures of line segment ab when the two legs are 24 and 7 using pythagorean theorem

OpenStudy (jiteshmeghwal9):

(7)^2+(x)^2=(24)^2 49+x^2=576 x^2=576-49 x^2=527 now, u have to find the sqrt of 527 for the value of x